# End-to-End Fine-Tuning of 3D Texture Generation using Differentiable Rewards

### v3

@@ Abstract
While recent 3D generative models can produce high-quality texture images, they often fail to capture human preferences or meet task-specific requirements. Moreover, a core challenge in the 3D texture generation domain is that most existing approaches rely on repeated calls to 2D text-to-image generative models, which lack an inherent understanding of the 3D structure of the input 3D mesh object. To alleviate these issues, we propose an end-to-end differentiable, reinforcement-learning-free framework that embeds human feedback, expressed as differentiable reward functions, directly into the 3D texture synthesis pipeline. By back-propagating preference signals through both geometric and appearance modules of the proposed framework, our method generates textures that respect the 3D geometry structure and align with desired criteria. To demonstrate its versatility, we introduce three novel geometry-aware reward functions, which offer a more controllable and interpretable pathway for creating high-quality 3D content from natural language. By conducting qualitative, quantitative, and user-preference evaluations against state-of-the-art methods, we demonstrate that our proposed strategy consistently outperforms existing approaches. We will make our implementation code publicly available upon acceptance of the paper.
{strip}[Uncaptioned image]
Figure 1: Texture images generated before and after reward fine‑tuning on different 3D mesh objects using different rewards. Each column shows the text prompt at the top and the corresponding reward function used for texture generation at the bottom. Texture images produced by our method consistently outperform the pre‑fine‑tuning baseline (InTeX [30]) across all rewards and experiments.
@@ 1 Introduction
While large-scale generative computer vision models learn broad knowledge and some reasoning skills to generate images [3, 11], videos [32], and 3D objects [23, 16], achieving precise control of their behavior is difficult due to the completely unsupervised nature of their training. Among these generative models, 3D generative ones often rely on adapting text-to-image models at inference [23, 16, 34], but the resulting 3D content may not align with human preferences or task-specific needs. This challenge highlights the need for a preference learning frameworks adapted to 3D content creation. One general solution proposed mainly in the large language models’ literature [20, 6, 9, 25] is to incorporate human or automated feedback as a supervisory signal to guide the generative models toward desired behavior. This is usually done by leveraging reinforcement learning from human feedback (RLHF) [6]. Prior work has shown that RLHF can significantly improve results in some domains, such as text generation [20, 9] and image synthesis [3, 11]. However, to the best of our knowledge, DreamReward [33] is the only prior work that incorporates human feedback into 3D generative model training. It applies differentiable rewards to optimize NeRF-based geometry editing, without training the generative model. In contrast, we directly fine-tune the diffusion model for texture editing within a 3D generation pipeline, enabling precise control aligned with user preferences. Moreover, unlike DreamReward’s aesthetic rewards that ignore geometry, our geometry-aware rewards ensure textures are both perceptually rich and structurally aligned with the geometry of the input 3D mesh. Despite recent advances, a core challenge in 3D texture generation remains: most existing approaches rely on repeated calls to 2D text-to-image models, which lack an inherent understanding of 3D structure. Therefore, applying preference learning directly to these 2D components often leads to solutions that neglect crucial 3D constraints. To address this, we propose an end-to-end differentiable preference learning framework that back-propagates differentiable rewards through the entire 3D generative pipeline, making the process inherently geometry-aware. By directly fine-tuning the 3D texture generation model through differentiable geometry-aware rewards, our framework supports interactive and user-aligned texture editing, producing results that are both visually compelling and geometrically coherent. The main contributions of this study are:
@@ 4 Experiments and Results
To show the effectiveness of our approach, we perform four training experiments using the four proposed geomatry-aware reward functions in Sec. 3.3, each detailed in the following. We qualitatively and quantitatively compare our results against InTeX [30] and TEXTure [26], two well-established state-of-the-art methods in the field of texture generation. We report the qualitative results for each experiment in Fig. 3, and Fig. 4, and quantitative results in Tab. 1 and Tab. 2 (more results in Appendix H).
Quantitative Comparison. As shown at the top of Tab. 1, we report the reward values for both our post-fine-tuned method and the pre-fine-tuned baseline (InTeX [30]) across five experiments. Each experiment is repeated five times, and we report the mean and standard deviation of the resulting reward values. Our method consistently outperforms InTeX numerically. In the following, we detail each experiment and provide a qualitative comparison against two strong baselines: TEXTure [26] and InTeX [30].
@@
Texrure Features Emphasis. This task aims to learn texture images with salient features (e.g., edges) emphasized at regions of high surface bending, represented by the magnitude of mean curvature. This promotes texture patterns that highlight 3D surface structure while preserving perceptual richness through color variation. Fig. 4 (right) shows qualitative results on a rabbit mesh with brick patterns. As illustrated, our method enhances texture features, such as edges and mortar, in proportion to local curvature, a capability InTex [30] lacks, often resulting in pattern-less (white) areas, particularly on the back and head of the rabbit (more results in Fig. 12).
Figure 4: Left: Qualitative results of the camera-pose-aware texture generation experiment on three different 3D mesh objects (female head, dragon, and male head). Our method produces textures with more aesthetic patterns that are more realistic and align more closely with the mesh’s semantic parts. Middle: Qualitative results for the symmetry‑aware experiment on a 3D balloon mesh. For each viewpoint, we present the rendered mesh with a vertical dashed line indicating the symmetry axis. Compared to InTeX [30] and TEXTure [26], our method produces textures that exhibit consistent patterns across symmetric regions of the mesh, confirming the reward’s ability to enforce symmetry. In contrast, without symmetry supervision, the textures on opposite sides often diverge significantly. Right: Qualitative results of the texture features emphasis experiment on a rabbit (bunny) object. We show the rendered 3D bunny from multiple viewpoints. Our method enhances texture features, such as edges and mortar, in proportion to mean curvature, a capability InTeX[30] and TEXTure [26] lacks, often resulting in pattern-less (white) areas, particularly on the back and head of the rabbit.
@@
Symmetry-Aware Texture Generation. The goal of this experiment is to encourages texture consistency across symmetric regions of a 3D object (see Algorithm 1). We evaluate this by training our pipeline using the symmetry reward on a balloon mesh object. Fig. 4 (middle) presents qualitative results for the experiments with symmetry-aware reward. For each viewpoint, we present the rendered mesh with a vertical dashed line indicating the symmetry axis. Compared to InTeX [30] and TEXTure [26], our method produces textures that exhibit consistent patterns across symmetric regions of the mesh, confirming the reward’s ability to enforce symmetry. In contrast, without symmetry supervision, the textures on opposite sides often diverge significantly (more results in Fig. 9).
@@
Camera-Pose-Aware Texture Generation. The objective of this experiment is to learn optimal camera viewpoints such that, when the object is rendered and textured from these views, the resulting texture maximizes the average aesthetic reward. Consequently, by maximizing this reward, the model becomes invariant to the initial camera positions: regardless of where the cameras start, the training will adjust their azimuth and elevation to surround the 3D object in a way that yields high‑quality, aesthetically pleasing textures. Fig. 4 (left) presents qualitative results on three different meshes, a female head, a miniature dragon, and a male head, comparing our fine‑tuning approach against InTeX [30] and TEXTure [26]. All methods share identical initial camera parameters; once initialized, each algorithm proceeds to paint the object. As shown in Fig. 4, our method consistently outperforms both baselines, producing high‑fidelity, geometry‑aware textures across all shapes. This robustness stems from the aesthetic‑reward signal guiding camera adjustment during texture reward learning. In contrast, InTeX [30] and TEXTure [26] remain sensitive to poorly chosen initial viewpoints and thus often fail to generate coherent, high‑quality textures under suboptimal camera configurations. More results on this experiment are provided in Fig. 8.
@@ Acknowledgements
We acknowledge funding from the Mila Tech Transfer Grant with Silicolabs and internal funds Concordia University. Computational resources were provided by Calcul Québec and the Digital Research Alliance of Canada.
@@ Appendix A Remarks on The Texture Generation Step
where $\vec{V}$ is the view vector, a vector starting from the camera position points toward a point on the surface of the 3D mesh object, $\vec{N}$ is the surface normal, vector perpendicular to the mesh surface, and $\phi$ is the angle between the view vector and surface normal. During the rendering process, this helps determine how to project and render 3D objects onto a 2D screen, ensuring that objects are viewed from the correct perspective. More specifically, the view direction cosine is computed for each rendered viewpoint, and the rendered viewpoint is partitioned into three regions based on the value of view direction cosine: 1) generate - the regions have been viewed for the first time by the camera and have never been painted during the texturing process, 2) refine - the regions that have already been viewed and painted from a viewpoint in previous iterations, but is now seen from a better camera angle (i.e. higher view direction cosine values) and should be painted again, and 3) keep - the regions that have already been viewed and pained from a good camera angle and should not be painted again. Sec. A.1 presents further details and the differentiable mathematical representation of our texture generation step.
@@ A.1 Mathematical Representation
where $f_{controlnet}$ represents the ControlNet model whose parameters are fixed during the training, $h^{i}$ is the differentiable renderer output, $y^{i}_{t-1}$ is the control signal generated by the ControlNet model at timestep $t$ for viewpoint $i$, and $c$ is the prompt embdedding obtained from the CLIP encoder model [24]. To enforce the consistency during the painting process, we follow [2] to incorporate latent mask blending in the latent space of the diffusion process. Specifically, by referring to the part that we wish to modify as $m_{blended}$ and to the remaining part as $m_{keep}=1-m_{blended}$, we blend the two parts in the latent space, as the diffusion progresses. This modifies the sampling process in a way that the diffusion model does not change the keep regions of the rendered image. Therefore, the latent variable at the current diffusion timestep $t$ is computed as:
@@ F.2 Geometry-Aware Texture Colorization Experiment and Results
This task aims to colorize textures based on surface bending intensity, represented by mean curvature, an average of the minimal and maximal curvature directions, on the 3D mesh. Specifically, the model is encouraged to apply warm colors (e.g., red, yellow) in high-curvature regions and cool colors (e.g., blue, green) in low-curvature areas. Fig. 11 shows qualitative results on rabbit and cow mesh objects colorirzed with different textual prompts. As illustrated, our method consistently adapts texture colors, regardless of initial patterns, according to local curvature and successfully maps warmth and coolness to geometric variation (more results in Fig. 11).
@@ Appendix G Attempts to Overcome Reward Hacking
Appendix H Additional Experiments and Results
We show expanded results from the main paper featuring more texture image and more viewpoints of different 3D objects. Fig. 9, Fig. 10, Fig. 11, and Fig. 12 shows the qualitative results of the symmetry-aware, geometry-texture alignment, geometry-guided texture colorization, and texture features emphasis experiments, respectively. Moreover, a quantitative comparison on the bunny model is presented in Tab. 2, using various texture patterns. It evaluates our fine-tuned method against InTeX [30] (before fine-tuning) for geometry–texture alignment. Our approach yields significantly higher alignment counts, demonstrating that the generated textures respect the underlying geometric structure, whereas the InTeX [30] fails to achieve such alignment.

### v4

@@ Abstract
While recent 3D generative models can produce high-quality texture images, they often fail to capture human preferences or meet task-specific requirements. Moreover, a core challenge in the 3D texture generation domain is that most existing approaches rely on repeated calls to 2D text-to-image generative models, which lack an inherent understanding of the 3D structure of the input 3D mesh object. To alleviate these issues, we propose an end-to-end differentiable, reinforcement-learning-free framework that embeds human feedback, expressed as differentiable reward functions, directly into the 3D texture synthesis pipeline. By back-propagating preference signals through both geometric and appearance modules of the proposed framework, our method generates textures that respect the 3D geometry structure and align with desired criteria. To demonstrate its versatility, we introduce three novel geometry-aware reward functions, which offer a more controllable and interpretable pathway for creating high-quality 3D content from natural language. By conducting qualitative, quantitative, and user-preference evaluations against state-of-the-art methods, we demonstrate that our proposed strategy consistently outperforms existing approaches. Our implementation code is publicly available at: https://github.com/AHHHZ975/Differentiable-Texture-Learning
@@ 1 Introduction
While large-scale generative computer vision models learn broad knowledge and some reasoning skills to generate images [4, 12], videos [36], and 3D objects [25, 18], achieving precise control of their behavior is difficult due to the completely unsupervised nature of their training. Among these generative models, 3D generative ones often rely on adapting text-to-image models at inference [25, 18, 40], but the resulting 3D content may not align with human preferences or task-specific needs. This challenge highlights the need for a preference learning frameworks adapted to 3D content creation. One general solution proposed mainly in the large language models’ literature [22, 7, 10, 28, 34] is to incorporate human or automated feedback as a supervisory signal to guide the generative models toward desired behavior. This is usually done by leveraging reinforcement learning from human feedback (RLHF) [7]. Prior work has shown that RLHF can significantly improve results in some domains, such as text generation [22, 10] and image synthesis [4, 12, 8]. However, to the best of our knowledge, DreamReward [39] is the only prior work that incorporates human feedback into 3D generative model training. It applies differentiable rewards to optimize NeRF-based geometry editing, without training the generative model. In contrast, we directly fine-tune the diffusion model for texture editing within a 3D generation pipeline, enabling precise control aligned with user preferences. Moreover, unlike DreamReward’s aesthetic rewards that ignore geometry, our geometry-aware rewards ensure textures are both perceptually rich and structurally aligned with the geometry of the input 3D mesh. Despite recent advances, a core challenge in 3D texture generation remains: most existing approaches rely on repeated calls to 2D text-to-image models, which lack an inherent understanding of 3D structure. Therefore, applying preference learning directly to these 2D components often leads to solutions that neglect crucial 3D constraints. To address this, we propose an end-to-end differentiable preference learning framework that back-propagates differentiable rewards through the entire 3D generative pipeline, making the process inherently geometry-aware. By directly fine-tuning the 3D texture generation model through differentiable geometry-aware rewards, our framework supports interactive and user-aligned texture editing, producing results that are both visually compelling and geometrically coherent. The main contributions of this study are:
@@ 4 Experiments and Results
To show the effectiveness of our approach, we perform four training experiments using the four proposed geomatry-aware reward functions in Sec. 3.3, each detailed in the following. We qualitatively and quantitatively compare our results against InTeX [33], TEXTure [29], Text2Tex [5], and Paint3D [40], four well-established state-of-the-art methods in the field of texture generation. We report the qualitative results for each experiment in Fig. 3, and Fig. 9, and quantitative results in Tab. 1, Tab. 2, and Tab. 4 (more results in Appendix I).
Quantitative Comparison. We report a variety of widely-used automatic metrics to evaluate visual quality, image-text alignment and human preference proxies in Tab. 2. Moreover, as shown at the top of Tab. 1, we report the reward values for both our post-fine-tuned method and the pre-fine-tuned baseline (InTeX [33]) across five experiments. Each experiment is repeated five times, and we report the mean and standard deviation of the resulting reward values. For each scene we render the textured mesh from multiple viewpoints. In Tab. 2, for each method and scene we compute the selected scalar metric per rendered view, then average those per-view values to obtain a 3D-scene-level score for that method. Finally, we report the mean of those scene-level scores across scenes. As presented in Tab. 2, our method consistently outperforms InTeX [33], TEXTure [29], Text2Tex [5], and Paint3D [40] numerically. Below we briefly define each metric and give the intended interpretation (higher = better or lower = better). (i) Aesthetic (LAION predictor) [32]: an image-level aesthetic score produced by a learned aesthetic predictor trained using LAION-style data. It is intended to correlate with general photographic/artistic quality and larger values indicate more aesthetically pleasing images. (ii) ImageReward [38]: a learned reward model trained on human preferences that predicts which images humans prefer given a prompt or context. We use the ImageReward model as a proxy for human preference; larger scores indicate texture images more likely to be preferred by annotators. (iii) HPSv2 (Human Preference Score v2) [37]: an improved preference-model variant tuned to better correlate with human judgments on text-to-image outputs. Like other preference scorers, higher is better. (iv) PickScore [3]: a preference model trained on the Pick-a-Pic dataset [16] to predict which of two candidate images a human would choose. We report the model’s scalar output aggregated across views. Higher is better. (v) CLIPScore [27]: the cosine similarity between image and text embeddings from CLIP [26]. We compute per-view CLIP similarity between the render image and the prompt and average across views. This measures image–text alignment and higher values indicate better alignment. (vi) Inference time: average time to produce a final textured mesh (or to run the texture generation/inference procedure for one object) reported in seconds. Lower is better; all timing measurements were performed on a machine equipped with a single NVIDIA RTX 4090 GPU. In the following, we detail each experiment and provide a qualitative comparison against three strong baselines (see Fig. 9 in Appendix I): InTeX [33], TEXTure [29], Text2Tex [5], and Paint3D [40].
Table 2: Quantitative comparison of our method against baselines across multiple automatic metrics. We report higher-is-better ($\uparrow$) scores for Aesthetic [32], ImageReward [38], HPSv2 [37], PickScore [3], and CLIPScore [27], and a lower-is-better ($\downarrow$) measure for inference time. Across these metrics, our method consistently outperforms all four baselines.
Method Aesthetic $\uparrow$ ImageReward $\uparrow$ HPSv2 $\uparrow$ PickScore $\uparrow$ CLIPScore $\uparrow$ Inference Time (sec) $\downarrow$
TEXTure [29] 4.3910 $\pm$ 0.0328 -1.2245 $\pm$ 0.1199 0.1728 $\pm$ 0.0031 20.0719 $\pm$ 0.1083 0.3142 $\pm$ 0.0051 150
Text2Tex [5] 4.4454 $\pm$ 0.0387 -1.5088 $\pm$ 0.0823 0.1769 $\pm$ 0.0052 19.6886 $\pm$ 0.1862 0.2900 $\pm$ 0.0054 450
InTeX [33] 4.7467 $\pm$ 0.0381 -1.0859 $\pm$ 0.1469 0.1879 $\pm$ 0.0033 19.8832 $\pm$ 0.2368 0.3077 $\pm$ 0.0055 15
Paint3D [40] 4.7192 $\pm$ 0.0388 -1.6621 $\pm$ 0.0977 0.1549 $\pm$ 0.0089 18.8966 $\pm$ 0.2798 0.2798 $\pm$ 0.0065 30
Ours - Cam-Pose-Aware Reward 4.9328 $\pm$ 0.0271 -0.0479 $\pm$ 0.1748 0.2095 $\pm$ 0.0058 19.7609 $\pm$ 0.1573 0.3003 $\pm$ 0.0057 15
Ours - Geo-Tex-Align Reward 4.7722 $\pm$ 0.0324 -0.1758 $\pm$ 0.1136 0.2479 $\pm$ 0.0034 21.5022 $\pm$ 0.1293 0.3367 $\pm$ 0.0041 15
Ours - Sym-Aware Reward 4.9473 $\pm$ 0.0309 0.0063 $\pm$ 0.0901 0.2118 $\pm$ 0.0031 20.8361 $\pm$ 0.0444 0.3076 $\pm$ 0.0032 15
Ours - Tex-Emphasis Reward 5.0308 $\pm$ 0.0376 -0.5960 $\pm$ 0.1694 0.2534 $\pm$ 0.0027 20.8822 $\pm$ 0.1125 0.3158 $\pm$ 0.0028 15
@@
Texture Features Emphasis. This task aims to learn texture images with salient features (e.g., edges) emphasized at regions of high surface bending, represented by the magnitude of mean curvature. This promotes texture patterns that highlight 3D surface structure while preserving perceptual richness through color variation. Fig. 9 (right) shows qualitative results on a rabbit mesh with brick patterns. As illustrated, our method enhances texture features, such as edges and mortar, in proportion to local curvature, a capability all baselines lack, often resulting in pattern-less (white) areas, particularly on the back and head of the rabbit (more results in Fig. 14).
@@
Symmetry-Aware Texture Generation. The goal of this experiment is to encourages texture consistency across symmetric regions of a 3D object (see Algorithm 1). We evaluate this by training our pipeline using the symmetry reward on a balloon mesh object. Fig. 9 (middle) presents qualitative results for the experiments with symmetry-aware reward. For each viewpoint, we present the rendered mesh with a vertical dashed line indicating the symmetry axis. Compared to the baselines, our method produces textures that exhibit consistent patterns across symmetric regions of the mesh, confirming the reward’s ability to enforce symmetry. In contrast, without symmetry supervision, the textures on opposite sides often diverge significantly (more results in Fig. 11).
@@
Camera-Pose-Aware Texture Generation. The objective of this experiment is to learn optimal camera viewpoints such that, when the object is rendered and textured from these views, the resulting texture maximizes the average aesthetic reward. Consequently, by maximizing this reward, the model becomes invariant to the initial camera positions: regardless of where the cameras start, the training will adjust their azimuth and elevation to surround the 3D object in a way that yields high‑quality, aesthetically pleasing textures. Fig. 9 (left) presents qualitative results on three different meshes, a female head, a miniature dragon, and a male head, comparing our fine‑tuning approach against baselines. All methods share identical initial camera parameters; once initialized, each algorithm proceeds to paint the object. As shown in Fig. 9, our method consistently outperforms all baselines, producing high‑fidelity, geometry‑aware textures across all shapes. This robustness stems from the aesthetic‑reward signal guiding camera adjustment during texture reward learning. In contrast, InTeX [33], TEXTure [29], and Paint3D [40] remain sensitive to poorly chosen initial viewpoints and thus often fail to generate coherent, high‑quality textures under suboptimal camera configurations. More results on this experiment are provided in Fig. 10.
@@ Acknowledgements
We acknowledge funding from the Mila Tech Transfer Grant with Silicolabs and internal funds from Concordia University. Computational resources were provided by Calcul Québec and the Digital Research Alliance of Canada.
@@ Appendix A Remarks on The Texture Generation Step
where $\vec{V}$ is the view vector, a vector starting from the camera position points toward a point on the surface of the 3D mesh object, $\vec{N}$ is the surface normal, vector perpendicular to the mesh surface, and $\phi$ is the angle between the view vector and surface normal. During the rendering process, this helps determine how to project and render 3D objects onto a 2D screen, ensuring that objects are viewed from the correct perspective. More specifically, the view direction cosine is computed for each rendered viewpoint, and the rendered viewpoint is partitioned into three regions based on the value of view direction cosine: 1) generate - the regions have been viewed for the first time by the camera and have never been painted during the texturing process, 2) refine - the regions that have already been viewed and painted from a viewpoint in previous iterations, but is now seen from a better camera angle (i.e. higher view direction cosine values) and should be painted again, and 3) keep - the regions that have already been viewed and painted from a good camera angle and should not be painted again. Sec. A.1 presents further details and the differentiable mathematical representation of our texture generation step.
@@ A.1 Mathematical Representation
where $f_{controlnet}$ represents the ControlNet model whose parameters are fixed during the training, $h^{i}$ is the differentiable renderer output, $y^{i}_{t-1}$ is the control signal generated by the ControlNet model at timestep $t$ for viewpoint $i$, and $c$ is the prompt embedding obtained from the CLIP encoder model [26]. To enforce the consistency during the painting process, we follow [2] to incorporate latent mask blending in the latent space of the diffusion process. Specifically, by referring to the part that we wish to modify as $m_{blended}$ and to the remaining part as $m_{keep}=1-m_{blended}$, we blend the two parts in the latent space, as the diffusion progresses. This modifies the sampling process in a way that the diffusion model does not change the keep regions of the rendered image. Therefore, the latent variable at the current diffusion timestep $t$ is computed as:
@@ F.2 Geometry-Aware Texture Colorization Experiment and Results
This task aims to colorize textures based on surface bending intensity, represented by mean curvature, an average of the minimal and maximal curvature directions, on the 3D mesh. Specifically, the model is encouraged to apply warm colors (e.g., red, yellow) in high-curvature regions and cool colors (e.g., blue, green) in low-curvature areas. Fig. 13 shows qualitative results on rabbit and cow mesh objects colorized with different textual prompts. As illustrated, our method consistently adapts texture colors, regardless of initial patterns, according to local curvature and successfully maps warmth and coolness to geometric variation (more results in Fig. 13).
@@ Appendix G Attempts to Overcome Reward Hacking
Appendix H Ablation Studies
We include a focused ablation study to quantify the effect of regularization choices, LoRA scaling, and the memory/computation trade-off of our proposed configuration. Results and qualitative examples are shown in Fig. 8 (regularization ablations) and Fig. 7 (LoRA scaling ablations), and a compact memory/throughput comparison is reported in Tab. 3.
Full vs. LoRA+Checkpointing fine-tuning.
Tab. 3 summarizes the memory and throughput trade-offs. Using LoRA in combination with activation checkpointing reduces the peak GPU memory substantially (from an infeasible full-fine-tuning footprint to a modest working set) and reduces the number of trainable parameters by several orders of magnitude, while maintaining acceptable throughput. The table demonstrates that our default configuration provides a realistic balance between memory cost and training performance for high-resolution texture adaptation. Therefore, when training on limited GPU memory, it is recommended to enable activation checkpointing and LoRA to dramatically reduce peak memory, as shown in Tab. 3.
Fine-tuning with/without regularization strategies.
Fig. 8 compares two regularizers that we found important in practice: a colorfulness regularizer (Fig. 8(a)) and a LoRA-based regularization term that penalizes LoRA parameters changes (Fig. 8(b)). Each subfigure shows three rows for the same example mesh: the input (before fine-tuning), the output after fine-tuning without the regularizer, and the output after fine-tuning with the regularizer. The colorfulness regularizer reduces undesirable saturations and preserves a more balanced color distribution across the texture image, which is visible in the bunny texture (Fig. 8(a)). The LoRA regularization term helps keep the fine-tuned model’s distribution close to the pre-trained one while allowing parameter updates to satisfy the reward objective. This yields texture images that both exhibit natural patterns and meet specific objectives (e.g., symmetry patterns in Fig. 8(b). Qualitatively, both regularizers improve perceptual texture quality compared to training without them, though the more beneficial choice depends on the reward and desired visual characteristics.
LoRA scaling.
Fig. 7 reports reward curves for multiple LoRA scale values under two reward objectives: symmetry-aware (Fig. 7(a)) and aesthetic (Fig. 7(b)). The plots show smoothed rewards over training iterations for each LoRA scale value. Across both objectives we observe a trade-off: larger LoRA scales accelerate early gains (they provide more expressiveness and thus faster improvement on the target reward), but they can also increase variance and, in some settings, cause less stable long-term behaviour. Smaller scales train more conservatively and tend to be more stable, but require more iterations to reach the same reward level. The two rewards behave slightly differently: the symmetry reward benefits noticeably from higher early expressiveness, while the aesthetic reward shows a more gradual, steady improvement for moderate scales. Based on these results, we recommend a moderate-to-high default (e.g., 0.7–0.9) and advise practitioners to tune upward only when faster short-term gains are needed and the risk of added variance is acceptable.
Table 3: Ablation study comparing LoRA+checkpointing fine-tuning with full fine-tuning. Results show a substantial reduction in memory usage and trainable parameters when using LoRA with activation checkpointing. (OOM = Out of Memory)
Configuration Peak GPU mem (GB) #Params (M) Throughput (sec/iter)
Full fine-tuning 300 (OOM) 859.5 OOM
Ours (LoRA + Checkpointing) 20 0.797 40
(a) Ablation on LoRA scale measured with the symmetry-aware reward. Curves show smoothed reward versus training iterations for multiple LoRA scale values. Larger LoRA scales produce faster initial gains on the symmetry objective but can introduce higher variance; moderate-to-high scales achieve a good balance between speed and stability.
(b) Ablation on LoRA scale measured with the aesthetic reward. Curves show the smoothed aesthetic reward over iterations for the same set of LoRA scales. The relative ranking of scales differs slightly from the symmetry objective, indicating that reward choice interacts with LoRA expressiveness.
Figure 7: Ablation study on different values of LoRA scale. (Top) symmetry-aware reward. (Bottom) aesthetic reward. In both plots each series corresponds to a different LoRA scale (see legends); curves were smoothed for visual clarity.
(a) Ablation on colorfulness regularization. Each row shows (top to bottom): (1) result after fine-tuning with colorfulness regularization, (2) result after fine-tuning without colorfulness regularization, and (3) the input / before fine-tuning baseline. Columns show a set of canonical views (side / back / side2 / front / top) and the unwrapped texture. The colorfulness regularizer reduces excessive saturation and preserves perceptually balanced color distributions in the texture image while keeping geometric detail.
(b) Ablation on LoRA scaling regularization. Rows are arranged as in Fig. 8(a): (1) after fine-tuning with LoRA regularization, (2) after fine-tuning without LoRA regularization, and (3) before fine-tuning. Columns show side, top and the unwrapped UV texture. Using a mild-to-high LoRA regularization term helps keep the fine-tuned model’s distribution close to the pre-trained one while allowing parameter updates to satisfy the reward objective. This yields texture images that both exhibit natural patterns and meet specific objectives (e.g. symmetry patterns here).
Figure 8: Ablation studies on regularization strategies. (Top) colorfulness regularization; (Bottom) LoRA sacling regularization. Each subfigure compares the same mesh and texture image before fine-tuning, after fine-tuning without the regularizer, and after fine-tuning with the regularizer (see subcaptions for details).
Appendix I Additional Experiments and Results